# Multi-orbital two-particle self-consistent approach – strengths and limitations

### v1

@@ II Multi-orbital TPSC
$1=3\neq 2=4\quad\quad\implies\quad\quad$ $\chi^{S^{1/2}S^{1/2}}_{o_{1},o_{1},o_{2},o_{2}}=2\left(\pm\braket{n_{o_{1}}^{\downarrow\uparrow}n_{o_{2}}^{\downarrow\uparrow}}-\braket{n_{o_{2}o_{1}}^{\downarrow}n_{o_{1}o_{2}}^{\uparrow}}\right),$ (36)
$1=4\neq 3=2\quad\quad\implies\quad\quad$ $\chi^{S^{1/2}S^{1/2}}_{o_{1},o_{2},o_{2},o_{1}}=\braket{n_{o_{2}}}+2\left(\mp\braket{n_{o_{1}}^{\downarrow\uparrow}n_{o_{2}}^{\downarrow\uparrow}}-\braket{n_{o_{2}}^{\downarrow}n_{o_{1}}^{\uparrow}}\right),$ (37)
$1=2\neq 3=4\quad\quad\implies\quad\quad$ $\chi^{S^{1/2}S^{1/2}}_{o_{1},o_{2},o_{1},o_{2}}=2\braket{n_{o_{1},o_{2}}^{\uparrow\downarrow}n_{o_{1},o_{2}}^{\downarrow\uparrow}}.$ (38)
@@
$1=3\neq 2=4\quad\quad\implies\quad\quad$ $\chi^{S^{3}S^{3}}_{o_{1},o_{1},o_{2},o_{2}}=2\braket{n_{o_{1}}^{\uparrow}n_{o_{2}}^{\uparrow}-n_{o_{1}}^{\uparrow}n_{o_{2}}^{\downarrow}}=-2\braket{n_{o_{2}o_{1}}^{\downarrow}n_{o_{1}o_{2}}^{\uparrow}},$ (39)
$1=4\neq 3=2\quad\quad\implies\quad\quad$ $\chi^{S^{3}S^{3}}_{o_{1},o_{2},o_{2},o_{1}}=2\braket{n_{o_{2}}^{\uparrow}(1-n_{o_{1}}^{\uparrow})}-2\braket{n_{o_{1}o_{2}}^{\uparrow}n_{o_{2}o_{1}}^{\downarrow}}=\braket{n_{o_{2}}}-2\braket{n_{o_{2}}^{\downarrow}n_{o_{1}}^{\uparrow}}.$ (40)
@@
In summary, the self-consistency loop consists of the Ansatz equations for TPSC3 (Eq. (9,10,11)) or TPSC5 (Eq. (12,13,14)), from which the vertex follows in the form of Eq. (24) or Eq. (23). The vertex in turn determines the susceptibility, calculated according to Eq. (25), from which the TPEVs can be extracted utilizing the sum rules (Eq. (41,42,43,44)). This loop can be solved in various ways – in its core it is a minimization or multidimensional root finding problem. Once a minimum is found, the TPEVs are used to determine the charge vertex by fitting to the sum rules in the charge sector. These are derived analogously to the spin sum rules and read for pair-wise identical indices:
$1=3\neq 2=4\quad\quad\implies\quad\quad$ $\chi_{n_{o_{1}o_{1}}n_{o_{2}o_{2}}}=2\braket{n_{o_{1}}^{\uparrow}n_{o_{2}}^{\uparrow}+n_{o_{1}}^{\uparrow}n_{o_{2}}^{\downarrow}}-\braket{n_{o_{1}}}\braket{n_{o_{2}}},$ (45)
$1=4\neq 3=2\quad\quad\implies\quad\quad$ $\chi_{n_{o_{1}o_{2}}n_{o_{2}o_{1}}}=2\braket{n_{o_{2}}^{\uparrow}(1-n_{o_{1}}^{\uparrow})}+2\braket{n_{o_{1}o_{2}}^{\uparrow}n_{o_{2}o_{1}}^{\downarrow}}-\braket{n_{o_{1}o_{2}}}\braket{n_{o_{2}o_{1}}},$ (46)
$1=2\neq 3=4\quad\quad\implies\quad\quad$ $\chi_{n_{o_{1}o_{2}}n_{o_{1}o_{2}}}=-2\braket{n_{o_{1}o_{2}}^{\uparrow\downarrow}n_{o_{1}o_{2}}^{\downarrow\uparrow}}-\braket{n_{o_{1}o_{2}}}\braket{n_{o_{1}o_{2}}}.$ (47)
@@
II.1 Differences to earlier approaches
Multi-orbital TPSC formalisms were already presented in Ref. [87] and Ref. [88]. Here, we briefly discuss the differences between the proposed approach and these earlier attempts of generalizing TPSC to a multi-orbital setting. First and foremost, the proposed procedure here is fully self-consistent, i.e. no additional symmetry constraints have to be enforced beyond the structure of the bare vertex. The self-consistent double occupancies seem to cure the issue of negative components of the charge vertex - thus the internal consistency check is fulfilled as long as both minimizations converge to a solution. However, utilizing more sum rules complicates the numerical root search, so that the computational cost of the present approach is higher. We include both the particle-hole symmetrized and the usual version of the Ansätze in our implementation. However, in this paper, we only consider half-filled systems, so that the results between the particle-hole symmetry enforcing and the usual Ansatz do not differ. Apart from these differences, the sum rules utilized in Ref. [87] and Ref. [88] are a subset of the sum rules employed here, and we checked that by constraining the equations we can reproduce the previous results.
@@ II.1 Differences to earlier approaches
III Benchmarking
Before benchmarking the approaches numerically, we will try to provide an analytical understanding of when the approach is suitable and when it it not. The model utilized in these benchmarks are variations of a multi-orbital Hubbard model, where each site contains two orbitals, see Fig. 1. If not mentioned otherwise, the two orbitals are not kinetically coupled and we set $t=1$ and give all other quantities relative to $t$. Furthermore, the orbitals have a coupling on the two-particle level due to the Hubbard-Kanamori interaction, which is parameterized by the on-site interaction $U$ and the Hunds coupling $J$. Such a model and its variations have been studied extensively in former works [88, 19, 93], to which we will compare our results. To analyze this model analytically, we consider two extreme cases, the one in which no Hunds coupling is present and the case in which the Hunds-coupling is $U/3$.
III.1 Model with pure density-density interactions
First, let us consider the case of a multi-orbital model with density-density interactions only. For the kinetic part of the Hamiltonian we consider a two-orbital square-lattice model without hopping between the orbitals,
$H_{kin}=\sum_{\braket{i,j},o_{1},o_{2},s}tc^{\dagger}_{i,o_{1},s}c_{j,o_{2},s}\;,$ (48)
where $i,j$ label the lattice sites. We furthermore assume half-filling. Due to the absence of inter-orbital hoppings the non-interacting susceptibility has a block diagonal form:
$\chi^{0}_{o_{1},o_{4},o_{2},o_{3}}(q,\omega)=G_{o_{1}o_{4}}(k,\nu)G_{o_{2}o_{3}}(k-q,\nu-\omega)=\delta_{o_{1},o_{4}}\delta_{o_{2},o_{3}}\chi^{0}(q,\omega)\,,$ (49)
where $\chi^{0}(q,\omega)$ is the susceptibility of the single band model. In this model we assume a pure density-density type interaction (meaning zero Hund’s coupling $J$)
@@ III.1 Model with pure density-density interactions
This interaction in combination with the kinetic term leads to a vanishing $\Sigma^{0}$, so that we do not have to consider it in the following. Since $J=0$ there is no spin selectivity and we thus expect $\braket{n_{0}^{\uparrow}n_{1}^{\uparrow}}=\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}$. Further, since $U=U^{\prime}$ there is no energy difference in double occupying the same or different orbitals, and hence $\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}=\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$. This also implies that the double occupancy $\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$ should decrease less rapidly with $U$ than for $U^{\prime}=0$. With this in mind we now turn to TPSC.
@@
$\tilde{D}_{o_{4},o_{1}}$ $=D_{o_{4},o_{1}}\frac{\sum_{s_{4}}\bar{\delta}_{s_{1},s_{4}}^{o_{1},o_{4}}\braket{n_{o_{4}}^{s_{4}}n_{o_{1}}^{s_{1}}}}{\sum_{s_{4}}n_{o_{4}}^{s_{4}}n_{o_{1}}^{s_{1}}-n_{o_{1}o_{4}}^{s_{1}s_{4}}n_{o_{4}o_{1}}^{s_{4}s_{1}}}\,.$ (51)
As we are at half-filling and have no inter-orbital coupling the denominator simplifies to
$0.5-0.25\delta_{o_{1},o_{4}}\;.$ (52)
@@
$\chi_{o_{1},o_{2},o_{3},o_{4}}(q,\omega)$ $=\sum_{o_{1}^{\prime},o_{3}^{\prime}}\delta_{o_{1},o_{4}}\delta_{o_{2}^{\prime},o_{3}^{\prime}}\chi^{0}(q,\omega)\left(\frac{1}{\mathbb{I}_{a^{\prime}b^{\prime}}\delta_{a^{\prime},d^{\prime}}\delta_{b^{\prime},c^{\prime}}-\delta_{a^{\prime},d^{\prime}}\delta_{e^{\prime},f^{\prime}}\chi^{0}(q,\omega)\tilde{U}\delta_{e^{\prime},f^{\prime}}\delta_{b^{\prime},c^{\prime}}}\right)_{o_{1}^{\prime},o_{2},o_{3},o_{4}^{\prime}}$ (53)
$=\delta_{o_{1},o_{4}}\delta_{o_{2},o_{3}}\chi^{0}(q,\omega)\frac{1}{1-\chi^{0}(q,\omega)\tilde{U}}\,,$ (54)
@@
where the fraction in the first line denotes a matrix inversion.
Figure 2: Comparison between TPSC3 (blue), TPSC5 (yellow) and ED (red) for the model defined in Eq. (55) at $\beta=0.5$, for different $U$ and $J$ combinations. We include the zeroth order self-energy in both TPSC variants.
@@
We note that the interacting susceptibility is identical to the one of a single orbital model with the same interaction value. The only non-zero components of the susceptibility are $\chi_{1111}$ and $\chi_{1221}$. Thus, from Eq. (41) and Eq. (42) it immediately follows that $\braket{n_{0}^{\uparrow}n_{1}^{\uparrow}}=\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}=\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$, as we expected from the arguments above. In contrast to our expectations, the specific form of the susceptibility implies that the converged TPSC loop will result in the same double occupancy as the one for the one-orbital model. In other words, adding an inter-orbital interaction of the same strength as the bare interaction does not make a difference for the double occupation predicted by the method. This is in contrast to the physically expected result discussed above – in the single orbital model in the strong coupling limit, we would expect zero double occupancy. However in this model, different states in which no orbital is doubly occupied and states in which all orbitals are doubly occupied are degenerate in energy so the double occupancy should decay much slower (as it is seen in ED, see Fig. 3a).
@@
To pinpoint what TPSC is missing in this case, we consider two coupled dimers (2-site) for which we compare the results between ED and TPSC. The behavior described above is reproduced as expected, see Fig. 3a. Further, since the susceptibility does not gain any strong $\tau$ dependence, implying that the frequency dependence also does not change drastically, as shown in Fig. 3b, we expect the vertex to be still reasonably well described by the static limit. The questions therefore are, first, where does the deviation stem from and second, why does this issue occur here but not in the single orbital model?
Let us first answer the second question. In the single-orbital case we fix the local and static expectation value to be the exact one. Crucially, the single site contains no further internal structure meaning that the exact local expectation value is always proportional to the non-interacting one. Thus representing this expectation value locally by a renormalized Hartree-Fock expectation value works, i.e., we can always get the correct value of the double occupancy by multiplying the non-interacting double occupancy with a single number. Multi-orbital systems have additional degrees of freedom, which enter the local and static contribution. We still approximate the TPEVs in a Hartree-Fock fashion, thus as a single Slater-determinant which intrinsically is connected to the non-interacting limit. Importantly, the local Hilbert-space is now spanned by more than a single wave-function and all of these states contribute to the expectation value. Thus, introducing interactions can fundamentally alter the structure of the eigenstates in the local Hamiltonian, so that the expectation value can deviate from a behavior representable by the non-interacting one. As an illustration, let us considering a dimer at half-filling with a Hubbard-Kanamori interaction. At $t=0$ all $S=1$ states are degenerate producing vastly different TPEV’s. In TPSC only one of these three states is represented by the Hartree-Fock decoupling and the others are missed entirely (thus behaving like the single orbital model in the absence of $J$). Thereby the basic idea of TPSC, which relies on fixing the local and static expectation values, breaks down. Further, this immediately suggests that the issue arises due to the Hartree-Fock decoupling, which is central to TPSC and thus there is no fix to this problem without altering the structure of the method.
III.2 Analysis of the model with $J=U/3$
If $J=U/3$, the same-spin density-density interaction becomes exactly zero. This choice is interesting, as it is the limiting case where the approach in Ref. [88] and TPSC resulted in the best agreement compared to DMFT, as is shown in Fig. 4. What is the origin of this apparent improvement? First, we should note that this is not a magic value at which the approach suddenly works, but there appears to be a steady improvement with increasing $J$. Second, we observe that for $J=U/3$ the same-spin interaction is exactly 00 and stays at that value throughout the calculation in TPSC5. In TPSC3 this interaction will gain a nonzero value, leading to the instability we observe in Fig. 4.
Since we are again mainly interested in the quality of the TPEV’s predicted from the spin sum rules, for now let us consider the half-filled dimer, a single unit cell of the form sketched in Fig. 1. In this toy model, the spectrum consists of three groups; the $S=1$ group with eigenvalue $U-3J$ and the $S=0$ group which is split into a two-fold degenerate pair at $U-J$ and a single eigenvalue with $U+J$. Thus, while the point $U=3J$ is not special, the separation between the excited state and the ground state grows linearly with $J$. Furthermore, the $S=1$ subspace is kind of special, as two states only contribute to a single TPEV ($\braket{n^{\uparrow}_{0}n^{\uparrow}_{1}}$) and the other state essentially behaves like in the single orbital model. Therefore, one should be able to approximate the TPEV’s by a scaling in all components but the $\braket{n^{\uparrow}_{0}n^{\uparrow}_{1}}$ one, which is incorrectly estimated because of the missing degenerate states. This is exactly what is observed for both TPSC5 and in Ref. [88]. Thus we conclude that in the end this improvement is related to the structure and number of states determining the local and static expectation values.
From these observations, we can extract a guiding principle of when to expect TPSC to work in a multi-orbital setting and when not. If there is a gap in the local spectrum between the ground state and the excited states, TPSC should perform better than when the states are close in energy. Further, when analyzing the nature of the states, we can extract which TPEV is expected to deviate strongly and which not. This also indicates that this specific part of TPSC performs better at lower temperatures (since it is basically a ground state targeting approach) - however, the local and static approximation of the vertex becomes more and more inappropriate at lower temperatures which is why TPSC typically fails at low temperatures.
In summary, the regimes in which multi-orbital TPSC is guaranteed to work well are the weak coupling limit and a small region in which excited states do not contribute to the local TPEVs (controlled by temperature and the gap), while the vertex still is local and static. The former is accessible by inspection, while the latter one is not clear a priori and requires other numerical simulations to be fully gauged. Notably, the issue described above is partially resolved when starting from DMFT TPEVs [89, 93].
III.3 Numerical benchmarks

### v3

@@ 2 Multi-orbital TPSC
$1=3\neq 2=4\;\implies\;$ $\chi^{S^{1/2}S^{1/2}}_{o_{1},o_{1},o_{2},o_{2}}=2\left(\pm\braket{n_{o_{1}}^{\downarrow\uparrow}n_{o_{2}}^{\downarrow\uparrow}}-\braket{n_{o_{2}o_{1}}^{\downarrow}n_{o_{1}o_{2}}^{\uparrow}}\right),$ (34)
$1=4\neq 3=2\;\implies\;$ $\chi^{S^{1/2}S^{1/2}}_{o_{1},o_{2},o_{2},o_{1}}=\braket{n_{o_{2}}}+2\left(\mp\braket{n_{o_{1}}^{\downarrow\uparrow}n_{o_{2}}^{\downarrow\uparrow}}-\braket{n_{o_{2}}^{\downarrow}n_{o_{1}}^{\uparrow}}\right),$ (35)
$1=2\neq 3=4\;\implies\;$ $\chi^{S^{1/2}S^{1/2}}_{o_{1},o_{2},o_{1},o_{2}}=2\braket{n_{o_{1},o_{2}}^{\uparrow\downarrow}n_{o_{1},o_{2}}^{\downarrow\uparrow}}.$ (36)
@@
$1=3\neq 2=4\;\implies\;$ $\chi^{S^{3}S^{3}}_{o_{1},o_{1},o_{2},o_{2}}=2\braket{n_{o_{1}}^{\uparrow}n_{o_{2}}^{\uparrow}-n_{o_{1}}^{\uparrow}n_{o_{2}}^{\downarrow}}=-2\braket{n_{o_{2}o_{1}}^{\downarrow}n_{o_{1}o_{2}}^{\uparrow}},$ (37)
$1=4\neq 3=2\;\implies\;$ $\chi^{S^{3}S^{3}}_{o_{1},o_{2},o_{2},o_{1}}=2\braket{n_{o_{2}}^{\uparrow}(1-n_{o_{1}}^{\uparrow})}-2\braket{n_{o_{1}o_{2}}^{\uparrow}n_{o_{2}o_{1}}^{\downarrow}}$ (38)
$\qquad\;\;\qquad=\braket{n_{o_{2}}}-2\braket{n_{o_{2}}^{\downarrow}n_{o_{1}}^{\uparrow}}.$
@@
In summary, the self-consistency loop consists of the Ansatz equations for TPSC3 (Eq. (11,12,13)) or TPSC5 (Eq. (14,15,16)), from which the vertex follows in the form of Eq. (22) or Eq. (21). The vertex in turn determines the susceptibility, calculated according to Eq. (23), from which the TPEVs can be extracted utilizing the sum rules (Eq. (39,40,41,42)). This loop can be solved in various ways – in its core it is a minimization or multidimensional root finding problem. Once a minimum is found, the TPEVs are used to determine the charge vertex by fitting to the sum rules in the charge sector. These are derived analogously to the spin sum rules and read for pair-wise identical indices:
$1=3\neq 2=4\;\implies\;$ $\chi_{n_{o_{1}o_{1}}n_{o_{2}o_{2}}}=2\braket{n_{o_{1}}^{\uparrow}n_{o_{2}}^{\uparrow}+n_{o_{1}}^{\uparrow}n_{o_{2}}^{\downarrow}}-\braket{n_{o_{1}}}\braket{n_{o_{2}}},$ (43)
$1=4\neq 3=2\;\implies\;$ $\chi_{n_{o_{1}o_{2}}n_{o_{2}o_{1}}}=2\braket{n_{o_{2}}^{\uparrow}(1-n_{o_{1}}^{\uparrow})}+2\braket{n_{o_{1}o_{2}}^{\uparrow}n_{o_{2}o_{1}}^{\downarrow}}$ (44)
$\qquad\qquad\qquad\qquad\qquad-\braket{n_{o_{1}o_{2}}}\braket{n_{o_{2}o_{1}}},$
$1=2\neq 3=4\;\implies\;$ $\chi_{n_{o_{1}o_{2}}n_{o_{1}o_{2}}}=-2\braket{n_{o_{1}o_{2}}^{\uparrow\downarrow}n_{o_{1}o_{2}}^{\downarrow\uparrow}}-\braket{n_{o_{1}o_{2}}}\braket{n_{o_{1}o_{2}}}.$ (45)
@@
2.1 Differences to earlier approaches
Multi-orbital TPSC formalisms were already presented in Ref. [89] and Ref. [90]. Here, we briefly discuss the differences between the proposed approach and these earlier attempts of generalizing TPSC to a multi-orbital setting. First and foremost, the proposed procedure here is fully self-consistent, i.e. no additional symmetry constraints have to be enforced beyond the structure of the bare vertex. The self-consistent double occupancies seem to cure the issue of negative components of the charge vertex - thus the internal consistency check is fulfilled as long as both minimizations converge to a solution. However, utilizing more sum rules complicates the numerical root search, so that the computational cost of the present approach is higher. We include both the particle-hole symmetrized (see App. D) and the usual version of the Ansätze in our implementation. However, in this paper, we only consider half-filled systems, so that the results between the particle-hole symmetry enforcing and the usual Ansatz do not differ. Apart from these differences, the sum rules utilized in Ref. [89] and Ref. [90] are a subset of the sum-rules employed here, see App. B.1, and we checked that by constraining the equations we can reproduce the previous results.
@@ 2.1 Differences to earlier approaches
3 Benchmarking
Before benchmarking the variants numerically, we will try to provide an analytical understanding of when the approach is suitable and when it is not. The models utilized in these benchmarks are variations of a multi-orbital Hubbard model, where each site contains two orbitals, see Fig. 1. The Hamiltonian of this model reads
$\hat{H}=\sum_{s,\braket{i,j},o}tc^{\dagger}_{o,s,i}c_{o,s,j}+\hat{H}_{HK}\,,$ (46)
where $i,j$ are lattice sites and $o$ labels the orbitals at each site. Here we introduced a Hubbard-Kanamori interaction, which has the following form
$H^{HK}=U\hat{n}_{o_{1^{\prime}}}^{\uparrow}\hat{n}_{o_{1^{\prime}}}^{\downarrow}+\sum_{{o_{1^{\prime}}}\neq{o_{2^{\prime}}}}$ $\left((U-2J)\hat{n}_{o_{1^{\prime}}}^{\uparrow}\hat{n}_{o_{2^{\prime}}}^{\downarrow}+(U-3J)\hat{n}_{o_{1^{\prime}}}^{s^{\prime}}\hat{n}_{o_{2^{\prime}}}^{s^{\prime}}\right.$ (47)
$\quad\left.-J\hat{c}^{{\dagger}}_{o_{1^{\prime}},\uparrow}\hat{c}_{o_{1^{\prime}},\downarrow}\hat{c}^{{\dagger}}_{o_{2^{\prime}},\downarrow}\hat{c}_{o_{2^{\prime}},\uparrow}+J\hat{c}^{{\dagger}}_{o_{1^{\prime}},\uparrow}\hat{c}^{{\dagger}}_{o_{1^{\prime}},\downarrow}\hat{c}_{o_{2^{\prime}},\downarrow}\hat{c}_{o_{2^{\prime}},\uparrow}\right)\,.$
[…]
$D_{o_{1},o_{4}}$ $=U-(1-\delta_{o_{1},o_{4}})2J\;,$ (48)
$C_{o_{1},o_{3}}$ $=(1-\delta_{o_{1},o_{3}})J\;,$ (49)
$P_{o_{1},o_{3}}$ $=(1-\delta_{o_{1},o_{3}})J\;.$ (50)
If not mentioned otherwise, the two orbitals are not kinetically coupled, we set $t=1$ and give all other quantities relative to $t$. The orbitals have a coupling on the two-particle level due to the Hubbard-Kanamori interaction, which is parameterized by the on-site interaction $U$ and the Hunds coupling $J$. Such a model and its variations have been studied extensively in former works [90, 19, 95], to which we will compare our results. To analyze this model analytically, we consider two limiting cases, the one in which no Hunds coupling is present and the case in which the Hunds-coupling is $U/3$.
3.1 Model with pure density-density interactions
First, let us consider the multi-orbital model described above with density-density interactions only. Furthermore, we assume half-filling, i.e. $\braket{n_{o_{1},s}}=0.5$, such that we have one particle per orbital. Due to the absence of inter-orbital hoppings the non-interacting susceptibility has a block diagonal form:
$\chi^{0}_{o_{1},o_{2},o_{3},o_{4}}(\boldsymbol{q},\omega)=G_{o_{1}o_{4}}(\boldsymbol{k},\nu)G_{o_{2}o_{3}}(\boldsymbol{k}-\boldsymbol{q},\nu-\omega)=\delta_{o_{1},o_{4}}\delta_{o_{2},o_{3}}\chi^{0}(\boldsymbol{q},\omega)\,,$ (51)
where $\chi^{0}(\boldsymbol{q},\omega)$ is the susceptibility of the respective single-band model, $\boldsymbol{q}$ is the momentum and $\omega$ the Matsubara frequency. Furthermore, we consider a pure density-density type interaction with identical intra- ($U$) and inter-orbital ($U^{\prime}$) Hubbard interaction (zero Hund’s coupling $J$):
@@ 3.1 Model with pure density-density interactions
This interaction in combination with the kinetic term leads to a vanishing $\Sigma^{0}$ (see App. (A.2) Eq. 99 or Ref. [91]). Since $J=0$ there is no spin selectivity and we thus expect $\braket{n_{0}^{\uparrow}n_{1}^{\uparrow}}=\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}$. Furthermore, since $U=U^{\prime}$, there is no energy difference for double occupying the same or different orbitals, and hence $\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}=\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$. This also implies that the double occupancy $\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$ should decrease less rapidly with $U$ than in the case with $U^{\prime}=0$ (which then is equivalent to the single orbital Hubbard model). With this in mind we now turn to TPSC.
@@
$\tilde{D}_{o_{4},o_{1}}$ $=D_{o_{4},o_{1}}\frac{\sum_{s_{4}}\braket{n_{o_{4}}^{s_{4}}n_{o_{1}}^{s_{1}}}}{\sum_{s_{4}}n_{o_{4}}^{s_{4}}n_{o_{1}}^{s_{1}}-n_{o_{1}o_{4}}^{s_{1}s_{4}}n_{o_{4}o_{1}}^{s_{4}s_{1}}}\,.$ (53)
As we are at half-filling and have no inter-orbital coupling, the Hamiltonian is diagonal in orbital space, i.e. $\braket{c^{\dagger}_{o_{1},s_{1}}c_{o_{2},s_{2}}}=0.5\delta_{o_{1},o_{2}}\delta_{s_{1},s_{2}}$. Therefore the denominator simplifies to
$\sum_{s_{4}}n_{o_{4}}^{s_{4}}n_{o_{1}}^{s_{1}}-n_{o_{1}o_{4}}^{s_{1}s_{4}}n_{o_{4}o_{1}}^{s_{4}s_{1}}=0.5-0.25\delta_{o_{1},o_{4}}\;.$ (54)
@@
$\chi_{o_{1},o_{2},o_{3},o_{4}}(\boldsymbol{q},\omega)$ $=\sum_{o_{1^{\prime}},o_{3^{\prime}}}\delta_{o_{1},o_{3^{\prime}}}\delta_{o_{1^{\prime}},o_{3}}\chi^{0}(\boldsymbol{q},\omega)\left(\frac{1}{\hat{1}-\hat{\tilde{D}}\hat{\chi}^{0}(\boldsymbol{q},\omega)}\right)_{o_{1^{\prime}},o_{2},o_{3^{\prime}},o_{4}},$ (55)
@@
where the fraction denotes an inversion. The matrix we need to invert reads in index notation
$\left(\hat{1}-\hat{\tilde{D}}\hat{\chi}^{0}(\boldsymbol{q},\omega)\right)_{o_{1},o_{2},o_{3},o_{4}}=\delta_{o_{1},o_{3}}\delta_{o_{2},o_{4}}-\delta_{o_{1},o_{3}}\delta_{o_{2^{\prime}},o_{4^{\prime}}}D_{o_{1},o_{2^{\prime}}}\chi^{0}(\boldsymbol{q},\omega)\delta_{o_{2^{\prime}},o_{4}}\delta_{o_{2},o_{4^{\prime}}}\,,$ (56)
which has a block diagonal structure. Utilizing that the inverse matrix of a block diagonal matrix is again block diagonal, we arrive at
$\chi_{o_{1},o_{3},o_{2},o_{4}}(\boldsymbol{q},\omega)$ $=\delta_{o_{1},o_{4}}\delta_{o_{2},o_{3}}\chi^{0}(\boldsymbol{q},\omega)\frac{1}{1-\chi^{0}(\boldsymbol{q},\omega)\tilde{D}}\,.$ (57)
We note that the interacting susceptibility is identical to the one of a single orbital model with the same renormalized interaction value. The only non-zero components of the susceptibility are $\chi_{1111}$ and $\chi_{1221}$. Thus, from Eq. (39) and Eq. (40) it immediately follows that $\braket{n_{0}^{\uparrow}n_{1}^{\uparrow}}=\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}=\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$, as we expected from the arguments above. Inserting these findings into the Ansatz equations yields
$\tilde{D}_{o_{4},o_{1}}=U\frac{\braket{n_{o_{1}}^{\uparrow}n_{o_{4}}^{\uparrow}}(1-\delta_{o_{1},o_{4}})+\braket{n_{o_{1}}^{\uparrow}n_{o_{4}}^{\downarrow}}}{0.5-0.25\delta_{o_{1},o_{4}}}=4U\braket{n_{o_{1}}^{\uparrow}n_{o_{4}}^{\downarrow}}\,.$ (58)
Due to the symmetries we found, the right hand side is orbitally independent and gives the same equation for the renormalized interactions as the single orbital case [80, 82]. This result is also independent of what filling we initially chose and is a general feature of the considered model.
@@
Therefore, in contrast to our expectations, the specific form of the susceptibility implies that the converged TPSC loop will result in the same double occupancy as the one for the one-orbital model. In other words, adding an inter-orbital interaction of the same strength as the bare interaction does not make a difference for the double occupation predicted by the method. This is in contrast to the physically expected result discussed above – in the single orbital model in the strong coupling limit, we would expect zero double occupancy. However in two-orbital model, different states in which no orbital is doubly occupied and states in which one orbitals is doubly occupied are degenerate in energy, so the double occupancy should decay more slowly.
@@
To pinpoint what TPSC is missing in this case we go back one step and consider a molecular system. Here, multi-orbital TPSC is essentially amounting to a small correction with respect to the underlying Hartree-Fock treatment, meaning we essentially benchmark the validity of the Hartree-Fock decoupling. The simplest such system we can write down are two coupled dimers, visualized for a four dimer case in Fig. 1, without Hund’s coupling $J$. The hopping $t$ again is set to $1$ and $U$ is varied. We compare the results between ED and TPSC, or better said Hartree-Fock. The behavior described above is reproduced, see Fig. 2a. Further, since the susceptibility does not gain any strong $\tau$ dependence, implying that the frequency dependence also does not change drastically, as shown in Fig. 2b. Therefore, we expect the vertex to be still reasonably well described by the static limit. The questions therefore are, first, where does the deviation stem from and second, why does this issue occur here but not in the single-orbital model?